# AssumptionBox linecolor=blue, outerlinewidth=2pt, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=10pt, innerleftmargin=10pt, backgroundcolor=gray!20

### v1

@@
Abstract.
This paper presents a framework for evaluating fairness in recommender systems powered by Large Language Models (RecLLMs), addressing the need for a unified approach that spans various fairness dimensions including sensitivity to user attributes, intrinsic fairness, and discussions of fairness based on underlying benefits. In addition, our framework introduces counterfactual evaluations and integrates diverse user group considerations to enhance the discourse on fairness evaluation for RecLLMs.
Our key contributions include the development of a robust framework for fairness evaluation in LLM-based recommendations and a structured method to create informative user profiles from demographic data, historical user preferences, and recent interactions. We argue that the latter is essential for enhancing personalization in such systems, especially in temporal-driven scenarios. We demonstrate the utility of our framework through practical applications on two datasets, LastFM-1K and ML-1M. We conduct experiments on a subsample of 80 users from each dataset, testing and assessing the effectiveness of various prompt construction scenarios and in-context learning, comprising more than 50 scenarios. This results in more than 4000 recommendations (80 × 50 = 4000). Our study reveals that while there are no significant unfairness issues in scenarios involving sensitive attributes, some concerns remain. However, in terms of intrinsic fairness, which does not involve direct sensitivity, unfairness across demographic groups remains significant. The code and data used for this paper are available at: https://shorturl.at/awBFM.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’24: ACM Symposium on Neural Gaze Detection, June 03–05, 2024, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction and Related work.
Context. The rapid evolution of Recommender Systems powered by Large Language Models (RecLLMs), such as ChatGPT, has demonstrated their capability to deliver personalized content across various domains (Spurlock et al., 2024; Vats et al., 2024; Jiang et al., 2024; Zhang et al., 2023; Li et al., 2023b). However, the deployment of these models in sensitive environments raises concerns about potential biases, considering that systems are essentially trained on large, unregulated datasets from the large-scale Internet that may inherently favor certain user demographics.
Our focus is on the direct use of LLMs for generating recommendations based on user-inputted textual prompts. As shown in Figure 1, for instance, a user has listened to songs such as “Flume” by Bon Iver and “In The End” by Linkin Park in the past, and recently listened to the songs “Points of Authority” and “Runaway” by Linkin Park. Based on this historical preference and recent interest, the user asks for the next top-3 suitable song recommendations after her/his final listening to “Aftiteling: Hold On To The World” by Kane. An illustration of the output recommendations and our evaluation notion is provided, showcasing examples of recommendation outputs when sensitive attributes are considered versus when they are not, or when they are presented in a counterfactual sense, each revealing potential discrepancies in fairness.
@@ Abstract.
Figure 1. This figure illustrates the direct use of Large Language Models (LLMs) in generating personalized recommendations. It compares outputs under neutral conditions with those generated under scenarios that consider sensitive attributes.
Related works. Fairness has long been recognized as a crucial issue in responsible and trustworthy recommender systems, garnering significant attention in recent years (Ekstrand et al., 2012; Li et al., 2023a; Deldjoo et al., 2024). As Large Language Models (LLMs) and generative models become increasingly prevalent, they introduce nuanced challenges in ensuring fairness. Recent research has started exploring the nuances of fairness in directly used LLMs, particularly when prompted with user inputs. Asn example, Zhang et al. (2023) propose a novel framework for formulating fairness in RecLLMs through the lens of sensitive attributes, suggesting that fairness entails similarity in recommendation lists in the absence and presence of these attributes. Extending this discussion, the authors in (Deldjoo and Di Noia, 2024) associate this notion to the concept of “recommendation benefits,” highlighting that different recommendations are not necessarily negative, and could be an indication of personalization. Other works have also begun examining this area (Vats et al., 2024). Our current work advances these aspects by better formalizing the problem (clear formalism for quantifying fairness), introducing novel concepts (e.g., counterfactual fairness), and exploring new tasks (sequential tasks). Our work also sheds light on better approaches for defining user profiles from the massive consumption profiles of users, which in our dataset can range from hundreds to thousands.
Research Questions and Contributions. This paper tackles critical research gaps in fairness evaluation for recommender systems powered by large language models (RecLLMs). We propose a novel, comprehensive framework named FairEvalLLM. It leverages existing fairness metrics while introducing new ones to provide a holistic assessment. Our contributions include:
  • Introduction and Formalization of an Evaluation Framework for RecLLMs: FairEvalLLM incorporates three notions of fairness specific to RecLLMs and tests each with two different underlying benefits (hit and rank). In particular, we have refined the conceptualization of fairness by discussing it in terms of underlying benefits and developing metrics specifically tailored to address the challenges presented by RecLLMs. Our framework focuses its discussion primarily on the consumer side as the main stakeholder. Moreover, we explore fairness not only within the real domain but also within a “counterfactual” domain, enriching our discussion of consumer fairness with scenarios that consider hypothetical alterations of sensitive attributes.
  • Profile Contribution Strategy: Another key novelty of our framework is the proposal for “”profile contribution strategy,” which involves creating detailed user profiles to simulate diverse demographic scenarios, considering the limitations of LLMs in terms of tokens. This strategy enables us to test the robustness of RecLLMs against various user inputs and uncover subtle biases that may not be evident through conventional methods. By systematically varying demographic and interaction components in these profiles, our work contributes to a better understanding of the impact of user modeling on assessing and mitigating potential biases within RecLLMs.
In light of the above contributions, we introduce the following practical research questions:
  1. RQ1.
    Are there major fairness issues arising from the use of RecLLMs in scenarios involving sensitive attributes such as gender and age, and how do different strategies (prompt construction, sampling, and ICL learning styles) affect, mitigate, or improve fairness in the recommendation outcomes?
  2. RQ2.
    How does intrinsic fairness manifest in RecLLMs across attribute-unaware scenarios, and which user groups benefit or suffer from potential biases in these settings?
2. Evaluation Framework for RecLLM Fairness
@@
We present FairEvalLLM, a multi-faceted framework designed to evaluate fairness in Recommender Systems powered by Large Language Models (RecLLMs). This work builds upon and extends significantly previous research (Zhang et al., 2023; Jiang et al., 2024; Deldjoo and Di Noia, 2024), introducing novel concepts that integrate and harmonize diverse perspectives on fairness.
2.1. Definitions.
2.1.1. Context
On the consumer side, our evaluation framework acknowledges a multifaceted understanding of fairness by considering various aspects: (i) the potential for discrimination when sensitive attributes are disclosed in prompts used by directly-employed LLMs; (ii) the implications of counterfactual scenarios where sensitive attributes are hypothetically altered; (iii) the neutrality of recommendations when sensitive attributes are absent. While (i) and (ii) are more specific to RecLLMs, noting that mainstream recommender systems do not typically use demographic information, the latter aspect focuses on assessing fairness across various user groups in comparison to a predefined ‘target representation’ (Kırnap et al., 2021; Amigó et al., 2023), aligning more closely with mainstream considerations.
We examine these three aspects in the fairness notion section (cf. Section 2.1.3); before that, we describe the rankers used to formalize/define the fairness notions under consideration, and what exactly constitutes the ‘reference’ ranking list, based on which fairness is gauged.
@@ 2.1.1. Context
For simplicity, we consider groups that are independent and binary. Overlapping groups and scenarios that consider the non-binary groups are beyond the scope of this paper and are left for future investigation.
2.1.2. Ranking Lists
We first introduce primary ranker types for fairness definitions.
Neutral Ranker ($\mathcal{R}_{N}$): Referred to as the neutral ranking list, this term describes a sequence of items $\{i_{1},i_{2},\ldots,i_{k}\}$ ranked by a Recommender Language Learning Model (RecLLM), using prompts that do not incorporate sensitive user attributes. The neutral ranker is designed to reflect scenarios based purely on non-sensitive demographic data, ensuring that the selection of recommended products is free from biases related to sensitive attributes such as gender, age, or other demographic characteristics. It bases recommendations solely on the historical interaction of the user with the system.
@@ 2.1.2. Ranking Lists
2.1.3. Fairness Notions
On the consumer side, we consider the following fairness notions, each linked to the corresponding rankers:
Neutral vs. Sensitive Ranker Deviation (NSD)::
@@ 2.1.3. Fairness Notions
Neutral vs. Counterfactual Sensitive Deviation (NCSD)::
@@
2.2. Fairness quantification
To quantify unfairness in Recommender Systems powered by Large Language Models (RecLLMs), we start by defining the general concept of benefit deviation, which serves as the foundation to quantify unfairness in our framework, given by:
(1) $\Delta\mathcal{B}=\mathcal{B}(\mathcal{R}_{X})-\mathcal{B}(\mathcal{R}_{ref})$
where $\mathcal{R}_{X}$ represent ranking generated under varying conditions, $\mathcal{R}_{ref}$ is the reference ranker (e.g., $\mathcal{R}_{N}$ in NSD), and $\mathcal{B}$ represents the benefit derived from each list. A higher value of $\Delta\mathcal{B}$ in every scenario below indicates a higher amount of unfairness.
  1. (1)
@@ 2.2. Fairness quantification
    (2) $\Delta\mathcal{B}=\mathcal{B}(\mathcal{R}_{S}^{a})-\mathcal{B}(\mathcal{R}_{N})$
@@
Table 1. Fairness Evaluation Based on the threshold $\delta$, $\Delta\mathcal{B}<\delta$ and p-value
@@
Level 1 Small - (p¿0.05) Safe
Level 2 Fairly large - (p¿0.05) Attention Needed
Level 3 Large - (p¿0.05) Likely Issue
Level 4 Large/Small - (p¡0.05) Significant Issue
Table 1 presents a structured assessment of fairness, organizing different levels of disparity based on $\Delta\mathcal{B}$ and associated p-values into categories ranging from ‘Safe’ to ‘Significant Issue’. This categorization helps stakeholders quickly identify potential biases in the recommendation system and determine the urgency of needed interventions. One might choose to adjust the number of levels or the criteria for each level based on their particular needs, regulations, and the nuances of their data.
@@
    (3) $\Delta\mathcal{B}=\mathcal{B}(\mathcal{R}_{CS}^{do(a)})-\mathcal{B}(\mathcal{R}_{N})$
@@
Finally, Intrinsic Fairness (IF) examines the fairness of recommendation distributions by a neutral ranker, $\mathcal{R}_{N}$, across sensitive groups such as male versus female. While the previous approaches may be more specific to RecLLMs due to the integration of “demographic information,” IF represents a more general approach that can also be and has been widely applied to traditional recommendation models, such as collaborative filtering models (Boratto et al., 2022; Ekstrand et al., 2012; Li et al., 2021). Essentially, IF evaluates whether the outcomes provided by $\mathcal{R}_{N}$ are fair by comparing the actual distribution of recommendations to a target (uniform) distribution across different demographic groups.
2.2.1. Benefit types.

### v2

@@
Abstract
The rapid adoption of large language models (LLMs) in recommender systems (RS) presents new challenges in understanding and evaluating their biases, which can result in unfairness or the amplification of stereotypes. Traditional fairness evaluations in RS primarily focus on collaborative filtering (CF) settings, which may not fully capture the complexities of LLMs, as these models often inherit biases from large, unregulated data. This paper proposes a normative framework to benchmark consumer fairness in LLM-powered recommender systems (RecLLMs). We critically examine how fairness norms in classical RS fall short in addressing the challenges posed by LLMs. We argue that this gap can lead to arbitrary conclusions about fairness, and we propose a more structured, formal approach to evaluate fairness in such systems. Our experiments on the MovieLens dataset on consumer fairness, using in-context learning (zero-shot vs. few-shot) reveal fairness deviations in age-based recommendations, particularly when additional contextual examples are introduced (ICL-2). Statistical significance tests confirm that these deviations are not random, highlighting the need for robust evaluation methods. While this work offers a preliminary discussion on a proposed normative framework, our hope is that it could provide a formal, principled approach for auditing and mitigating bias in RecLLMs. The code and dataset used for this work will be shared at gihub-anonymized.
keywords:
[…]
1 Introduction
Context. Fairness in recommender systems (RS) has garnered significant attention in recent years, driven by the need to mitigate biases that can negatively impact both consumers and providers. Most existing research in the RS field, however, has focused on fairness in classical collaborative filtering (CF) setting, which primarily rely on in-domain user-item interaction data to compute recommendations [1, 2]. Since the introduction of ChatGPT in 2023, the RS community has seen unprecedented interest in integrating generative models—particularly those powered by pre-trained large language models (LLMs)–for personalization [3, 4]. These models, which capture vast amounts of knowledge during pre-training on semi-supervised tasks, can be quickly adapted to a variety of contexts within RS, offering notable advantages for personalized recommendations. These benefits include efficiency (rapid deployment and adaptability), precision and context-awareness (enhanced personalization across diverse tasks), and robustness in data scarce scenarios (the ability to perform well under sparse data conditions).
However, integrating LLMs into recommender systems introduces new risks, particularly biases embedded in the training data, which can lead to unfairness or the amplification of stereotypes affecting sensitive or protected groups. Given the unregulated nature of online data, it becomes a critical concern to (i) understand, (ii) evaluate, and (iii) mitigate biases and unfairness of RecLLMs. RecLLMs differ significantly from traditional CF systems in several key areas: the input space (where simple star ratings are replaced by more complex inputs like natural language user profiles), model types (pre-trained on vast datasets rather than being directly trained using in-domain data), and output spaces (offering more structured outputs like complementary items or detailed explanations instead of just item IDs). This work advocates for developing RecLLM fairness evaluation frameworks that account for these differences, as they can influence our understanding of what is fair and unfair. The central question explored in this work revolves around the following:
How can we audit the fairness of RecLLMs (recommender systems powered by large language models), and how does fairness evaluation for RecLLMs differ from traditional CF methods?
Research Problem. In RecLLMs, there is the ability to incorporate sensitive demographic information, such as gender, directly from natural language (NL) user profiles—something that mainstream collaborative filtering (CF) models typically do not utilize. Figure 1 illustrates this with an example. Building on this, Zhang et al. [5] propose a fairness evaluation framework for RecLLMs that examines how demographic attributes, such as gender, can impact recommender outcomes. Their approach defines unfairness as the difference in recommendations between a sensitive ranker (which considers demographic factors) and a neutral ranker (which does not), based solely on differences in item IDs or their ranking in the list. This approach equates differences across user groups with unfairness, oversimplifying the issue by failing to account for situations where such differences might represent valid personalization.
For example, if a RecLLM recommends songs like “Hey Young Girl” by Lloyd based on prior interactions in a neutral scenario, but then suggests a song by Jamiroquai when gender is factored in, the system might flag this as unfair. This assumes the recommendation change is driven by gender stereotypes, without considering that such recommendations may align with the user’s actual preferences.
To address this limitation,  Deldjoo and Di Noia [6] propose the CFairLLM framework, which improves on FaiRLLM by assessing fairness based on whether the benefits a sensitive ranker provides differ (or more precisely worse) from those of a reference ranker, to flag it unfair. Thus, CFairLLM evaluates recommendation variations by comparing them to the user’s true preferences. For instance, if the sensitive ranker suggests songs that better match the user’s preferences, even if they differ from the neutral ranker’s list, these variations may indicate proper personalization rather than unfairness. This highlights the importance of defining fairness norms clearly; otherwise, research results might be contradictory.
Overall, these approaches lack a formal discussion of key elements essential for determining fairness in RecLLMs, such as establishing a clear reference point for evaluation and defining precise benefits and metrics for measuring fairness.
@@ 1 Introduction
Figure 1: This figure illustrates the direct use of Large Language Models (LLMs) in generating personalized recommendations. It compares outputs under neutral conditions with those generated under scenarios that consider sensitive attributes.
Contributions. The current work advocates for the need for a “normative framework” to address the above issues–one that ensures fairness is evaluated according to well-established, principled standards, rather than subjective or arbitrary assumptions (e.g., assuming all differences indicate unfairness). Additionally, it highlights the importance of having clear, objective criteria for consistently and meaningfully assessing fairness.
The main contributions of this paper are as follows:
  1. 1.
    Distinguishing Fairness Frameworks: We differentiate between two types of fairness evaluation in RecLLMs: (i) fairness when sensitive attributes are involved in generating recommendations (referred to as sensitive rankers), and (ii) fairness when comparing recommendations from neutral rankers, which do not use sensitive attributes, to predefined target distributions. This distinction is crucial for understanding the nuances in fairness evaluation between RecLLMs and traditional CF models, which typically rely on (i). Our work considers both approaches.
  2. 2.
    Introduction of Novel Fairness Metrics: Building on the previous distinction, we introduce three fairness evaluation metrics, which essentially operate on the same principle: comparing the deviation and difference between the RS ranker and another ranker (either a reference or a target representation). The metrics are as follows: Neutral vs. Sensitive Ranker Deviation (NSD), Neutral vs. Counterfactual Sensitive Deviation (NCSD), and Intrinsic Fairness (IF). While NSD and NCSD are specific to RecLLMs, IF is typically applied in CF scenarios.
  3. 3.
    Introducing of Reference Rankers: A key aspect of fairness evaluation in RecLLMs is the use of reference rankers. Each fairness notion relies on a specific reference ranker, such as the neutral ranker (for NSD and NCSD) or a predefined target distribution (for IF). The choice of the reference ranker plays a critical role in measuring deviations, as it influences how fairness or unfairness is perceived. In contrast to previous work that focuses merely on differences between rankers, as suggested by [6], we consider preference alignment with ground truth data.
  4. 4.
    Quantification of Fairness Deviations: To quantify ranker deviation, we propose metrics that assess the quality of recommendations through both set-based and rank-based measures. These measures examine the accuracy of the ranking and the overall benefit derived by different demographic groups. In addition to evaluating benefit deviation, we apply statistical significance tests to determine whether observed differences in recommendation benefits across groups are meaningful. This provides a comprehensive mechanism to quantify fairness deviations and ensures that the quality of recommendations is assessed based on rigorous, statistically sound methods.
Overall, this work advocates the need for a normative framework for fairness evaluation of RecLLMs, proposing a more formal approach than previous research. In this framework, fairness is evaluated against clear, well-defined standards, aiming to avoid arbitrary assumptions and provide a structured method for assessing how sensitive information affect recommendation outcomes. Our focus in this work is specifically on consumer fairness.
2 Evaluation Framework for Consumer Fairness in RecLLMs
@@
We present a multi-faceted framework designed to evaluate fairness in Recommender Systems powered by Large Language Models (RecLLMs). This work builds significantly upon and extends previous research [5, 6], but offers a more formal approach to fairness evaluation in RecLLMs.
2.1 Definitions.
@@
For simplicity, in this work we only consider groups that are independent and binary. However, overlapping groups can also be considered by defining combinations of attributes, such that the set of possible overlapping groups is given by $\mathcal{G}=\{(a_{1},b_{1}),(a_{1},b_{2}),(a_{2},b_{1}),(a_{2},b_{2})\}$. Here, each pair represents a distinct demographic group (e.g., $(a_{1},b_{1})$ for young males, $(a_{2},b_{2})$ for older females), allowing to analyze fairness across intersectional identities [6].
Ranking Lists. We first introduce primary ranker types for fairness definitions.
Neutral Ranker ($\mathcal{R}_{N}$): Referred to as the neutral ranking list, this term describes a sequence of items $\{i_{1},i_{2},\ldots,i_{k}\}$ ranked by a Recommender Language Learning Model (RecLLM), using NL profile (as prompts) that do not incorporate sensitive user attributes. The neutral ranker is designed to reflect scenarios based purely on non-sensitive demographic data. It bases recommendations solely on the historical interaction of the user with the system.
@@
Fairness Frameworks. On the consumer side, we consider the following fairness notions, each linked to the corresponding rankers:
Neutral vs. Sensitive Ranker Deviation (NSD):
@@
Neutral vs. Counterfactual Sensitive Deviation (NCSD):
@@
Fairness quantification. To quantify unfairness in RecLLMs, we start by defining the general concept of benefit deviation, which serves as the foundation to quantify unfairness in our framework, given by:
$\Delta\mathcal{B}=\mathcal{B}(\mathcal{R}_{X})-\mathcal{B}(\mathcal{R}_{ref})$ (1)
where $\mathcal{R}_{X}$ represent ranking generated by the target recommender (e.g., sensitive ranker), $\mathcal{R}_{ref}$ is the reference ranker (e.g., $\mathcal{R}_{N}$ in NSD), and $\mathcal{B}$ represents the benefit derived from each list. A lower value of $\Delta\mathcal{B}$ in every scenario below indicates a higher amount of unfairness.
  1. 1.
@@
    $\Delta\mathcal{B}=\mathcal{B}(\mathcal{R}_{S}^{a})-\mathcal{B}(\mathcal{R}_{N})$ (2)
@@
Table 1: Fairness Evaluation Based on the threshold $\delta$, $\Delta\mathcal{B}<\delta$ and p-value
@@
Level 1 Small - (p>0.05) Safe
Level 2 Fairly large - (p>0.05) Attention Needed
Level 3 Large - (p>0.05) Likely Issue
Level 4 Large/Small - (p<0.05) Significant Issue
Table 1 essentially aims to present a structured assessment of fairness, organizing different levels of disparity based on $\Delta\mathcal{B}$ and associated p-values into categories ranging from ‘Safe’ to ‘Significant Issue’. This categorization helps stakeholders quickly identify potential biases in the recommendation system and determine the urgency of needed interventions. One might choose to adjust the number of levels or the criteria for each level based on their particular needs, regulations, and the nuances of their data.
@@
    $\Delta\mathcal{B}=\mathcal{B}(\mathcal{R}_{CS}^{do(a)})-\mathcal{B}(\mathcal{R}_{N})$ (3)
@@
Finally, Intrinsic Fairness (IF) examines the fairness of recommendation distributions by a neutral ranker, $\mathcal{R}_{N}$, across sensitive groups such as male versus female. While the previous approaches may be more specific to RecLLMs due to the integration of “demographic information,” IF represents a more general approach that can also be and has been widely applied to traditional recommendation models, such as collaborative filtering models [7, 8, 9]. Essentially, IF evaluates whether the outcomes provided by $\mathcal{R}_{N}$ are fair by comparing the actual distribution of recommendations to a target (uniform) distribution across different demographic groups.
2.1.1 Benefit types.